# Bourbaki degree of pairs of projective surfaces

### 31 Mar 2026 v3

@@ Abstract
The logarithmic tangent sheaf associated to an algebraically independent sequence of homogeneous polynomials — defined as the kernel of the associated Jacobian matrix — naturally generalizes the classical logarithmic tangent sheaf of a divisor in a projective space to the case of subvarieties defined by more than one equation. As is the case for divisors, one may investigate the freeness of such sequences, and other weaker notions.
The present work focuses on sequences of two homogeneous polynomials in four variables. We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis. The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation of the degrees. We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants. Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension one foliation in $\mathbb{P}^{3}$.
We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic polynomials, relating stability and Chern classes with the discrete invariants introduced, while classifying free and nearly free cases. In particular, one of the nearly free examples induces an unstable, non-split tangent sheaf for a codimension one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
@@ 1 Introduction
Let $R=\kappa[x_{0},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$ variables with coefficients in an algebraically closed field $\kappa$. For an algebraically independent sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials in $R$ with degrees $d_{1}+1,\ldots,d_{k}+1$, respectively, with $d_{1}\leq\ldots\leq d_{k}$ one can consider the Jacobian matrix as a map of locally free sheaves on $\mathbb{P}^{n}\doteq\operatorname{Proj}R$
@@
The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the tangent logarithmic sheaf associated to the sequence $\sigma$. This definition is in analogy (see [undefj]) with the case of the tangent logarithmic sheaf associated to divisors in $\mathbb{P}^{n}$, namely for $k=1$. A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a sum of line bundles on $\mathbb{P}^{n}$.
In the case of divisors on $\mathbb{P}^{n}$, where $k=1$, one has a short exact sequence of the form:
@@
where $J_{f}=\langle\partial_{0}f,\ldots,\partial_{n}f\rangle\trianglelefteq R$ is the Jacobian ideal of the homogeneous polynomial $f$. The sheaf ${\mathcal{T}}_{f}$ is the sheaf of $\mathcal{O}_{\mathbb{P}^{n}}$-modules associated to the graded $R$-module $\operatorname{Syz}(J_{f})$ of Jacobian syzygies of $f$. For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated to the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
For the generalization of $k>1$, one should consider sequences rather than the associated ideals $I=(f_{1},\ldots,f_{k})$, as pointed out by [undefj, Example 2.7]. When $d_{1}=\ldots=d_{k}$, elements in sequences $\sigma$, $\sigma^{\prime}$ generating the same ideal vary only by an invertible constant matrix, and furthermore ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\sigma^{\prime}}$ (see [undefj, Lemma 2.14]).
We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a non-zero syzygy for the matrix $\nabla\sigma$, called the initial degree. Following the terminology of [undefj], we call a sequence $\sigma=(f_{1},\ldots,f_{k})$ compressible if, after a linear change of coordinates, there is a variable that does not occur in any of $f_{1},\ldots,f_{k}$. This is equivalent to $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=0$. Moreover, the number of variables which are independent give trivial copies ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{n}}^{m}\oplus\mathcal{E}$, where $\mathcal{E}$ is a logarithmic sheaf associated to the sequence $\sigma$ in the ring $\kappa[x_{0},\ldots,x_{n-m}]$ (see [undefj, Lemma 2.8]).
One may also consider the relationship with Bourbaki ideals, as in [undefr] and [undeff]. For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
@@
where $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection. The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice. In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [undefr]. By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. The focus of this work is to understand the concept above for $n=3$ and $k=2$.
@@
For $k=1$, the singular scheme of $V(f)$, defined by the Jacobian ideal $\Sigma_{f}\doteq V(J_{f})\subset\mathbb{P}^{n}$, plays an important role in this study. If $s=\dim\Sigma_{f}$, the Hilbert polynomial is given by
$H(\mathcal{O}_{J_{f}}(d),t)=H(R/J_{f},t)=\frac{\deg(\Sigma)}{s!}t^{s}+O(t^{s-1}),$
@@
and, by definition, the degree $\deg(\Sigma)$ is the leading coefficient above. Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$. For example, if $n=2$, $\deg(\Sigma)$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$, and it appears in the formula of the Bourbaki degree of a projective curve (given in [undefr]):
$\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}),$
where $e=\operatorname{indeg}({\mathcal{T}}_{f})$.
Although this is a natural definition for $n=2$, for $n>2$ one has to choose a set of generators for the logarithmic sheaf ${\mathcal{T}}_{f}$ to get an ideal sheaf as the cokernel. This problem is addressed in a recent work [undeff] where the authors consider another kind of Bourbaki sequences, recently developed in [undefq], by choosing generators of maximal degree.
@@
We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel of the Jacobian matrix, which corresponds to the coherent sheaf $\mathcal{O}_{J_{f}}(d)$ in the case $k=1$, supported at the singular scheme of $V(f)$. In [undefj], the authors introduce the Jacobian scheme of $\sigma$, denoted by $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$, as the zero locus of the $(k\times k)$-minors of the matrix $\nabla\sigma$. In particular, the reduced support of the sheaf $|\operatorname{supp}\mathcal{Q}_{\sigma}|$ coincides with the reduced support of $\Xi_{\sigma}$. The generic case is when $\dim(\mathcal{Q}_{\sigma})\leq n-2$, and in particular $c_{1}(\mathcal{Q}_{\sigma})=0$. In this case, we call $\sigma$ a normal sequence.
@@
For now on, we assume $n=3$ and $k=2$, that is, sequences $\sigma=(f,g)$ of homogeneous polynomials $f,g\in R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, and we set $d\doteq d_{f}+d_{g}$. We also assume sequences $\sigma$ are normal, as defined above, so that $c_{1}(\mathcal{Q}_{\sigma})=0$ and $c_{1}({\mathcal{T}}_{\sigma})=-d$. Since the logarithmic sheaf ${\mathcal{T}}_{\sigma}$ has rank two, compressibility implies freeness, and a normal sequence $\sigma$ is compressible if and only if ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)$.
If $\sigma$ is a normal sequence, $\dim(\mathcal{Q}_{\sigma})\leq 1$ and the Hilbert polynomial is of the form
$H(\mathcal{Q}_{\sigma},t)=\deg(\mathcal{Q}_{\sigma})t+O(1),$
@@
We set $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$, so $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ is zero-dimensional. From the associativity formula applied to the graded homogeneous module $\mathcal{Q}$ associated to $\mathcal{Q}_{\sigma}$, we obtain the expression (see, for example, [undefs, Theorem $14.7$])
$m(\sigma)=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(R/P)=2}\operatorname{{length}}_{R_{P}}\mathcal{Q}_{P}\cdot\deg(R/P).$
@@
Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we reproduce analogous considerations as in [undefr]. We show a formula for the Bourbaki degree of $\sigma$ given by
$\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma),$
so that $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. A first result of the paper is some bounds for the discrete quantities defined above.
@@
Theorem A(Theorem A) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials of the ring $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. Then:
@@
      $\sigma$ is compressible;
We also consider some other inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 2.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 8): for low enough values of $m(\sigma)$, the sequences are not free, and for high enough values of $\operatorname{Bour}(\sigma)$, the sequences must be $\mu$-stable. In general, the Bourbaki degree is bounded by $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, and the maximum is attained precisely by the matrices $\nabla\sigma$ for which the Buchsbaum–Rim complex is a minimal resolution, or equivalently $m(\sigma)=0$. Geometrically, all regular pencils ($d_{f}=d_{g}$) with isolated singularities satisfy this property (see Remark 9).
In Section 2.2, we relate our notion to the notion of the Bourbaki degree for a projective plane curve, by considering a reduced polynomial $g\in\kappa[x_{0},x_{1},x_{2}]$ together with its associated surface $V(g)\subset\mathbb{P}^{3}$ and projective curve $X=V(g)\subset\mathbb{P}^{2}$, to obtain $\operatorname{Bour}(X)=\operatorname{Bour}(g,x_{3})$.
Afterwards, in Section 2.3, we develop relationships between free resolutions for ${\mathcal{T}}_{\sigma}$ and for the ideal sheaf $\mathcal{I}_{B}$, where $B$ is the Bourbaki scheme obtained by the choice of a minimal syzygy. Related to the free resolutions, we introduce the notions of nearly free sequences and $3$-syzygy sequences, with the following chain of implications
$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3\text{-syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1.$
The converses do not hold, as we show by building examples with normal pencils of cubics $(d_{f}=d_{g}=2)$. We point out nearly free sequences are never locally free, as we show in 19.
In Section 3, we explore the structure of codimension one foliation of the sheaf ${\mathcal{T}}_{\sigma}(1)$, presented in [undefj, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$ using the sub-foliations by curves induced by these global sections. For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [undefr, Corollary 2.11]). The main theorem of the section is:
@@
In Section 4, we finish with a study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$. To do so, we use classical results on bounds for Chern classes of reflexive sheaves in $\mathbb{P}^{3}$. We classify all free and nearly free cases in these classes in terms of their discrete invariants and establish some stability results, which are summarized below:
Theorem C.(Theorem C)Let $\sigma=(f,g)$ be a normal pencil of cubic surfaces in $\mathbb{P}^{3}$. Then, if we denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
  • (b)
  • (c)
    There is only one case of nearly free sequence $\sigma$, with discrete invariants $m(\sigma)=7$, $e=2$ and $c_{3}({\mathcal{T}}_{\sigma})=2$ (see Example 42), which is strictly $\mu$-semistable;
  • (d)
In this case, we have an example of a strictly $\mu$-semistable logarithmic sheaf ${\mathcal{T}}_{\sigma}$ with $m(\sigma)=4$ (see Example 15), but the bound given above for stability may not be sharp.
Theorem D.(Theorem D) Let $\sigma=(f,g)$ be a normal sequence with $d_{f}=1,d_{g}=2$. Then, if $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$:
  • (a)
  • (b)
  • (c)
    There are two cases of nearly free sequences $\sigma$, both with $m(\sigma)=4$, one where ${\mathcal{T}}_{\sigma}$ is $\mu$-stable with $c_{3}({\mathcal{T}}_{\sigma})=1$ and another one where ${\mathcal{T}}_{\sigma}$ is $\mu$-unstable with $c_{3}({\mathcal{T}}_{\sigma})=3$ (see Example 56 and Example 14);
  • (d)
  • (e)
    If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Example 14 induces a codimension one foliation $\mathcal{F}$ of degree $3$ with tangent sheaf $T_{\mathcal{F}}={\mathcal{T}}_{\sigma}(1)$ which is non-split and not $\mu$-semistable. This example provides a negative answer to a conjecture posed by Calvo-Andrade, Correa and Jardim, in [undefa]:
Conjecture. If the tangent sheaf of a codimension one foliation on $\mathbb{P}^{3}$ is not split, then it is $\mu$-semistable.
Along the text, we describe examples developed computationally with aid of Macaulay2 software ([undefl]) for the two families of sequences.

### 27 Apr 2026 v4

@@ Abstract
The present work focuses on studying the logarithmic tangent sheaf associated with sequences of two homogeneous polynomials in four variables. We introduce two positive discrete invariants: the invariant $m$ and the Bourbaki degree of a sequence, inspired by the framework of the Bourbaki degree recently developed for projective plane curves by Jardim-Nejad-Simis. The invariant $m$ plays the role of the Tjurina number of plane projective curves and is bounded by a quadratic relation of the degrees. We establish results concerning the interplay of minimal degree for syzygies of the Jacobian matrix and the introduced discrete invariants. Our approach uses tools from foliation theory, taking advantage of the fact that the logarithmic sheaf is, up to a twist, the tangent sheaf of a codimension-one foliation in $\mathbb{P}^{3}$. We provide examples and classification results for pencils of cubics and for pairs of a quadric and a cubic. In particular, one of the nearly-free examples induces an unstable, non-split tangent sheaf for a codimension-one foliation of degree $3$, answering, in the negative, a conjecture of Calvo-Andrade, Correa and Jardim from $2018$.
@@ 1 Introduction
Let $\kappa$ be an algebraically closed field of characteristic zero and let $R=\kappa[x_{0},\ldots,x_{n}]$ be the polynomial ring in $n+1\geq 3$. For an algebraically independent sequence $\sigma=(f_{1},\ldots,f_{k})$ of homogeneous polynomials in $R$ with degrees $d_{f}+1,\ldots,d_{k}+1$, respectively, with $d_{f}\leq\ldots\leq d_{k}$ one can consider the Jacobian matrix as a map of locally free sheaves on $\mathbb{P}^{n}\doteq\operatorname{Proj}R$
@@
The kernel of $\nabla\sigma$, a reflexive sheaf of rank n+1kn+1-k on $\mathbb{P}^{n}$, is called the logarithmic tangent sheaf associated with the sequence $\sigma$. This definition is in analogy (see [11]) with the case of divisors in $\mathbb{P}^{n}$, namely for $k=1$. A sequence $\sigma$ is said to be free whenever the sheaf ${\mathcal{T}}_{\sigma}$ splits as a direct sum of line bundles on $\mathbb{P}^{n}$. In the case of divisors, one has a short exact sequence of the form:
@@
where $J_{f}=(\partial_{0}f,\ldots,\partial_{n}f)\subseteq R$ is the Jacobian ideal of the homogeneous polynomial $f$. The sheaf ${\mathcal{T}}_{f}$ is the sheaf of $\mathcal{O}_{\mathbb{P}^{n}}$-modules associated with the graded $R$-module $\operatorname{Syz}(J_{f})$ of Jacobian syzygies of $f$. For a sequence $\sigma$, the sheaf ${\mathcal{T}}_{\sigma}$ is associated with the graded $R$-module of syzygies of the Jacobian matrix $\nabla\sigma$.
We denote by $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$ the minimum degree for a nonzero syzygy of the matrix $\nabla\sigma$, called the initial degree. For k=n1k=n-1, a choice of global section of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ yields a short exact sequence
@@
where $B=B_{\nu}\subset\mathbb{P}^{n}$ is a pure codimension two scheme which is generically locally a complete intersection. The scheme $B_{\nu}$ depends on the choice of the syzygy $\nu$, but its Hilbert polynomial is independent of such a choice. In particular, we may consider the degree $\deg(B_{\nu})=\operatorname{Bour}(\sigma)$, which we call Bourbaki degree of the sequence $\sigma$, in analogy with the case studied for $k=1$ and $n=2$ in [17]. By construction, $\operatorname{Bour}(\sigma)=0$ if and only if $\sigma$ is a free sequence. The goal of this work is to study the concept above for $n=3$ and $k=2$.
@@
For $k=1$, the singular scheme $\Sigma_{f}$ of $V(f)$, defined by the Jacobian ideal, plays an important role. If $s=\dim\Sigma_{f}$, the Hilbert polynomial is given by
$H(\mathcal{O}_{J_{f}}(d),t)=H(R/J_{f},t)=\frac{\deg(\Sigma_{f})}{s!}t^{s}+O(t^{s-1}),$
@@
and, by definition, the leading coefficient $\deg(\Sigma_{f})$ is the degree of the singular scheme. Moreover, assuming the hypersurface $V(f)$ is reduced, we obtain $s\leq n-2$. For $n=2$, this means singularities are isolated, and $\deg(\Sigma_{f})$ coincides with the Tjurina number of the projective plane curve $V(f)\subset\mathbb{P}^{2}$. The formula of the Bourbaki degree in this case is ([17, Theorem 2.1]):
$\operatorname{Bour}(f)=e(e-d)+d^{2}-\deg(\Sigma_{f}).$
@@
We denote by $\mathcal{Q}_{\sigma}=\operatorname{coker}(\nabla\sigma)$ the cokernel of the Jacobian matrix and call the scheme $\Xi_{\sigma}\doteq V(\bigwedge^{k}\nabla\sigma)$ the Jacobian scheme of $\sigma$, given by the zero locus of the $(k\times k)$-minors of the matrix. The set-theoretical support of $\mathcal{Q}_{\sigma}$ is the reduced locus $(\Xi_{\sigma})_{\text{red}}$, although the scheme structure of both may differ. When $\dim(\mathcal{Q}_{\sigma})\leq n-2$, we call $\sigma$ a normal sequence.
@@
Hereafter, we assume $n=3$ and $k=2$, that is, sequences $\sigma=(f,g)$ of homogeneous polynomials $f,g\in R\doteq\kappa[x_{0},\ldots,x_{3}]$ with degrees $\deg(f)=d_{f}+1,\deg(g)=d_{g}+1$, and we set $d\doteq d_{f}+d_{g}$. We also assume sequences $\sigma$ are normal, as defined above, so $c_{1}(\mathcal{Q}_{\sigma})=0$ and $c_{1}({\mathcal{T}}_{\sigma})=-d$. Then $\dim(\mathcal{Q}_{\sigma})\leq 1$ and the Hilbert polynomial is of the form
$H(\mathcal{Q}_{\sigma},t)=m(\sigma)t+O(1),$
@@
where $m(\sigma)\doteq\deg(\mathcal{Q}_{\sigma})\geq 0$ and $m(\sigma)=0$ if and only if $\Xi_{\sigma}$ is zero-dimensional. From the associativity formula ([18, Theorem $14.7$]) applied to the graded homogeneous module $Q$ associated with $\mathcal{Q}_{\sigma}$, we obtain the expression
$m(\sigma)=\sum_{\mathfrak{p}\in\operatorname{\operatorname{Ann}}(\mathcal{Q}_{\sigma}),\dim(R/\mathfrak{p})=2}\operatorname{\operatorname{length}}_{R_{\mathfrak{p}}}Q_{\mathfrak{p}}\cdot\deg(R/\mathfrak{p}).$
@@
Using the fact that ${\mathcal{T}}_{\sigma}$ is a rank two reflexive sheaf, we reproduce analogous considerations as in [17]. We show the following formula for the Bourbaki degree of $\sigma$:
@@
Proposition: Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. If $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, then the Bourbaki degree of $\sigma$ is given by
$\operatorname{Bour}(\sigma)=e(e-d)+d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}-m(\sigma).$
A first result of the paper is some bounds for the discrete quantities defined above.
Theorem A(Theorem A) Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials in $R$, with degrees $d_{f}+1,d_{g}+1$ respectively, and set $d=d_{f}+d_{g}$. Then:
@@
      $\sigma$ is compressible, that is, up to a linear change of coordinates of $\mathbb{P}^{3}$, both polynomials in $\sigma$ are independent of some variable $x_{0},\ldots,x_{3}$;
We also show inequalities for $m(\sigma)$ and $\operatorname{Bour}(\sigma)$ in Section 2.1 related to freeness and $\mu$-stability of the sheaf ${\mathcal{T}}_{\sigma}$ (see 9). In general, the Bourbaki degree is bounded by $\operatorname{Bour}(\sigma)\leq d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}$, and the maximum is attained precisely whenever $m(\sigma)=0$. In this case, the cokernel $\mathcal{Q}_{\sigma}$ is minimally resolved by the Buchsbaum–Rim complex. In particular, ${\mathcal{T}}_{\sigma}(d)\simeq\operatorname{coker}(\nabla\sigma)^{T}$.
In Section 2.2, we consider sequences $\sigma=(x_{3},g)$, where $g\in\kappa[x_{0},x_{1},x_{2}]$ is a reduced homogeneous polynomial. We show that $\operatorname{Bour}(\sigma)$ coincides with the Bourbaki degree of the plane curve defined by $V(g)\subset\mathbb{P}^{2}=V(x_{3})$. In Section 2.3, we relate free resolutions of ${\mathcal{T}}_{\sigma}$ and of the ideal sheaf $\mathcal{I}_{B}$, where $B$ is the Bourbaki scheme obtained by the choice of a minimal degree syzygy. As a consequence, we introduce the notions of nearly free sequences and $3$-syzygy sequences, with the following chain of implications
$\sigma\text{ is nearly free}\Rightarrow\sigma\text{ is a }3\text{-syzygy sequence}\Rightarrow\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1,$
where $\operatorname{gpdim}({\mathcal{T}}_{\sigma})$ is the minimal length of a free resolution of the $R$-module $H^{0}_{*}({\mathcal{T}}_{\sigma})=\bigoplus_{l\in\mathbb{Z}}H^{0}({\mathcal{T}}_{\sigma}(l))$. The converses do not hold, see Example 16 and Example 23.
In Section 3, we explore the structure of codimension-one foliation associated with $\sigma$, presented in [11, Section 9], to obtain characterizations of low initial degrees $\operatorname{indeg}({\mathcal{T}}_{\sigma})\in\{1,2\}$. For plane projective curves, $e=1$ implies that $\operatorname{Bour}(f)\in\{0,1\}$ (see, for example, [17, Corollary 2.11]). The main theorem of the section is:
@@
We note that the bounds above are independent of the degrees $d_{f},d_{g}$. In Section 4, we study on two particular families of normal sequences: pencils of cubics $(d_{f}=d_{g}=2)$ and sequences with $d_{f}=1,d_{g}=2$, defining a degree $6$ curve inside a quadric surface in $\mathbb{P}^{3}$. To do so, we use classical results on bounds for Chern classes of reflexive sheaves in $\mathbb{P}^{3}$. We classify all free and nearly-free cases in these classes in terms of their discrete invariants and establish some stability results, which are summarized at the tables below (for precise statements, see Theorem Theorem C and Theorem Theorem D).
Finally, we provide a counter-example for a conjecture posed by Calvo-Andrade, Correa and Jardim, in [3]. They proposed that a tangent sheaf of a codimension-one foliation on $\mathbb{P}^{3}$ is either split or $\mu$-semistable. Example 15 gives a sequence $\sigma$ which induces a codimension-one foliation whose tangent sheaf is neither split nor $\mu$-semistable.
Along the text, we describe examples developed computationally with aid of Macaulay2 software ([13]) for the two families of sequences.
Classification results for $d_{f}=d_{g}=2$.
$m(\sigma)$ $\operatorname{Bour}(\sigma)$ Behavior minimal resolution for logarithmic tangent sheaf
12 0 free, compressible $\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-4)$
9 0 free, unstable $\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$
8 0 free, $\mu$-semistable $\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 2}$
7 1 unique nearly-free $0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0$
$\leq 6$ $\geq 2$ $\mu$-semistable
$\leq 2$ $\geq 6$ $\mu$-stable
0 12 $\mu$-stable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-6)^{\oplus 2}\xrightarrow{(\nabla\sigma)^{T}}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to{\mathcal{T}}_{\sigma}\to 0$
Classification results for $d_{f}=1$, $d_{g}=2$.
$m(\sigma)$ $\operatorname{Bour}(\sigma)$ Behavior minimal resolution for logarithmic tangent sheaf
7 0 free, compressible $\mathcal{O}_{\mathbb{P}^{3}}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)$
5 0 free, unstable $\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$
4 1 nearly-free, unstable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-3)\to\mathcal{O}_{\mathbb{P}^{3}}(-2)^{\oplus 3}\to{\mathcal{T}}_{\sigma}\to 0$
4 1 nearly-free, $\mu$-stable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\to\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\to{\mathcal{T}}_{\sigma}\to 0$
<3<3 $\geq 3$ $\mu$-stable
0 7 $\mu$-stable $0\to\mathcal{O}_{\mathbb{P}^{3}}(-4)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{(\nabla\sigma)^{T}}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 4}\to{\mathcal{T}}_{\sigma}\to 0$